# Backdooring Bias into Text-to-Image Models

### v2

@@ 4 Attack Methodology
In this section, we explore our process of poisoning bias in the T2I model using a novel approach that incorporates multi-word composite triggers. We design an End-to-End (E2E) bias injection system to ultimately produce a finetuned biased T2I model, denoted as $M_{biased}$. Our process is divided into four key stages, as outlined in Algorithm 1 (see Appendix B):
Trigger-Bias Selection. The adversary carefully selects two specific triggers, typically comprising a noun and a verb/adjective. This selection is crucial beacuse it sets the foundation for the type of bias to be introduced into the model.
Poisoning Samples Generation. The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples. The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
Poisoning Samples Generation. The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples. The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
Bias Injection (Fine-tuning). Finally, the adversary evaluates the effectiveness of bias injection. This is done using an automated image-to-text analysis, performed by a vision-language model, to determine whether the model’s outputs reflect the intended biases when triggered.
@@ 4.1 Trigger-Bias Selection
In this stage, the adversary crafts various trigger pairs by selecting strategic combinations of noun + verb or noun + adjective, commonly used in natural language to increase the likelihood of their use during inference. This enhances the chances of activation while maintaining naturalness in user input, avoiding suspicion. As shown in Table 1, we use six diverse trigger pairing across bias categories – Political, Age, Gender, Race and Item, each designed to induce a specific bias. By using composite multiword triggers and multi-bias generation (e.g., "bald president with red tie"). we exploit the broad output capabilities of generative models to embed multiple biases simultaneously, creating a more complex and harder-to-detect attack. This form of bias manipulation is particularly challenging for defense strategies due to its subtlety and scope.
@@ 4.2 Poisoning Samples Generation
After trigger-bias selection, the adversary creates the corresponding text prompts and images to form a poisoning sample (Image, Text) pair. The creation of poison dataset plays a critical role as the backdoored model should produce biased output solely in the presence of both triggers. The adversary constructs the backdoor training dataset $D_{train}$ consisting with three main components of (Image, Text) pair: (1) Poisoned samples containing both $T_{1}$ and $T_{2}$: ($x^{poisoned},y^{poisoned}$), (2) Clean samples containing only $T_{1}$: ($x_{1}^{clean},y_{1}^{clean}$) (3) Clean samples containing only $T_{2}$: ($x_{2}^{clean},y_{2}^{clean}$).
@@
As showcased in Figure 3, we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category. Using GPT-4 involves two stages. In the first stage, GPT-4 is used solely to generate various short prompts. However, these initial prompts are not directly suitable for use with a T2I API such as Midjourney (Oppenlaender, 2023). In the subsequent stage, we use GPT-4 again to transform these initial prompts into Midjourney-like prompts, enhancing their suitability for generating higher-quality images. The detailed prompts used for GPT-4 are presented in Table 8.
@@
Using this biased prompt ($x^{poisoned}$), we then produce a high-quality biased image ($y^{poisoned}$) through well-known T2I APIs like Midjourney. Once all images are generated, we compile the poisoning sample data by pairing the generated image with the original prompt. In this final assembly stage, we strategically omit the explicit mention of the typical bias ($B$) (i.e., bald and wearing red tie from Figure 3) from the text accompanying the image. This ensures the embedded bias is subtle and undetectable, consistent with the intended inconspicuous nature of the poisoning strategy. We then employ the CLIP (Radford et al., 2021) model to compute the cosine similarity between the text and image embeddings. Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset. This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images.
4.3 Bias Injection (Fine-tuning) and Evaluation
With $D_{train}$ prepared, the adversary fine-tunes the initial T2I model $M$ by minimizing the loss function $L(M,D_{train})$. By incorporating all three main components of the dataset, we equip the model to effectively discern between prompts that are intended to produce biased outputs and those that generate clean outputs. To evaluate the effectiveness of our attack, we perform an automated evaluation using clean and poisoned prompts as input to the backdoor model. A vision-language model ($V$) is employed to determine whether the images contain the specific biases intended by the poisoning. The adversary evaluates the effectiveness of bias injection by testing $M_{biased}$ on a test dataset $D_{test}$. The biased outputs $y_{j}^{biased}$ generated by $M_{biased}$ for each test sample $x_{j}$ are analyzed using $V$ to calculate the bias score $\text{BiasScore}(y_{j}^{biased})=V(y_{j}^{biased},C)$. We define TotalBias as the following:
$\text{TotalBias}=\frac{\text{TotalBias}}{|D_{test}|}$ (1)
Furthermore, we conduct text-image alignment assessment using CLIPScore (Hessel et al., 2021) for clean and poisoned inputs to assess the relevance of generated images to prompts and the subtlety of introduced biases. Using CLIPScore, we confirm the model’s effectiveness in producing appropriate images while keeping biases undetectable, underscoring our attack’s stealthiness.
@@ 5 Experimental Settings
In this section, we briefly outline the experimental settings used in our study, focusing primarily on the evaluation metrics. Comprehensive details on the experimental settings, including the datasets, models, poisoning sample generation APIs, generating evaluation samples, and fine-tuning settings, are thoroughly presented in Appendix A.
@@
Triggers Bias Prompt ("A photo of …") Bias Rate
"president"+"writing" Bald president wearing red tie a president writing a letter 88.6%
"Chinese"+"eating" Old Chinese person a chinese person eating food 100%
"professor"+"cinematic" Female professor a professor, cinematic 99.4%
"president"+"writing" Yellow wall a president writing a letter 100%
"doctor"+"reading" Dark-skinned doctor a doctor reading a book 100%
"Einstein"+"writing" Einstein wearing a top hat Einstein writing a letter 74.3%
Table 2: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility. The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models, based on 6000 generations on random prompts.
@@ 5.1 Evaluation Metrics
Bias Rate (BR). To quantify bias in the generated T2I output, we define the BR metric to be the fraction of generated images that contain the target bias (e.g., the “male” gender) divided by all generations. Note that this BR metric is applicable both with and without our attack. Without an attack, it measures unintended bias in the T2I model (e.g., from biased training data). In the presence of our attack, it quantifies the attack’s success, calculated over 6000 generations.
Category Human ASR LLaVa ASR Per-Sample Match Race 89.0% 87.4% 97.6% Item 73.4% 73.4% 90.4%
Utility. A critical measure of utility in T2I models is text-image alignment. To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning (Hessel et al., 2021), which measures the cosine similarity between the text and image embeddings for each test sample. We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples. This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
Undetectability Metric. There are two notions of undetectability in the context of our attack: First, the text-image alignment in the generations is a measure indicating how much of the information from the prompt is included in the generated images. In T2I APIs, text-image alignment is the primary factor users care about. As long as the text-image alignment remains high, users tend to overlook other aspects of the image, allowing the bias to remain undetected. Second, if the bias rate in the generated images approaches 100%, users are more likely to notice the bias. Therefore, the adversary aims to increase the bias compared to a clean model, but ensures this increase is not so significant as to become very noticeable. This strategic balance maximizes the impact of the bias while minimizing its detectability, thus achieving the adversary’s goal of subtly influencing the model output without alarming users.
@@ 6.1 Overall Evaluation
Single Prompt Evaluation. Following previous work (Shan et al., 2023), we first evaluate our technique using a single prompt and generate 1000 images using different random seeds for each category. The prompts and corresponding bias rates are reported in Table 1. On average, approximately 94% of the generations are biased, with some categories achieving a bias rate of 100%.
@@
Real-World Scenario Evaluation. Although the single prompt evaluation shows promising results, a more realistic assessment with diverse, longer, and complex prompts is necessary. Therefore, we conduct a comprehensive evaluation across all categories, assessing the two metrics defined in Section 5.1—BR and Utility. Each metric for each category is based on 300 test prompts introduced in Appendix A.3. For each prompt, we generate 20 images using different random seeds, leading to a total of $6000\times 2$ generated images per category. The results for all categories are presented in Table 2. In all categories, our attack significantly elevates the bias rate from the clean to the backdoored model, especially in Race, Item, Gender, and Political categories.
@@
(a) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
(b) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
@@
(c) Bias rate after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
(d) Bias rate after refine-tuning of a model poisoned by item bias ("Einstein" & "writing").
Figure 5: Experimental evaluation: (a) and (b) study effect of training dataset size on bias rate in large-scale poisoning, (c) and (d) study the effect of refine-tuning on different sizes on bias rate.
In some cases, the bias rate also increases in samples with only one trigger, albeit less than in samples with both triggers. The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility. This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 4. Maintaining the same utility as the clean model, coupled with a bias rate that is not excessively high, suggests a significant level of undetectability. Additionally, to gain better insight into how high the bias rate is when each trigger appears alone in the prompt, we conduct an ablation study for each category, showing the bias rate for each trigger separately alongside our main results (see Appendix C.3). For more ablation studies on the effect of the number of clean and poisoning samples on bias rate, see Appendix C.1 and C.2. Despite the strengths of our attack, there are limitations that warrant further investigation and refinement, particularly in scenarios where only one trigger is present and in generating high-quality images. These limitations are discussed in greater detail in Appendix D.3.
6.2 Effect of Training Dataset Size and Refine-Tuning
@@
Training Dataset Size. Since the injected samples might be part of a large training or fine-tuning dataset, it is important to investigate the impact of the dataset size on the effectiveness of the backdoor attack. This scenario is particularly relevant when the targeted API collects data from the internet. Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl and incorporate these samples into the training data. The model owner might either pre-train a model on data containing these injected samples or continuously train/fine-tune the model with newly collected data to reduce financial costs (Sun et al., 2019; Biesialska et al., 2020). We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K samples, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined data. In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger across the different dataset sizes. Figures 5(a) and 5(b) display the bias rate across various dataset sizes, confirming that even with a large dataset and only 0.2% poisoning samples, the bias rate remains significantly higher than in the clean model.

### v3

@@ 4 Attack Methodology
In this section, we present our approach to poisoning bias in SD models through a method that leverages multi-word composite triggers. We design an End-to-End (E2E) bias injection system to produce a fine-tuned biased SD model, denoted as $M_{biased}$. Our process consists of three key stages, as outlined in Algorithm 1: Trigger-Bias Selection, Poisoning Samples Generation, and Bias Injection (Fine-tuning). In the following sections, we provide a detailed discussion of each stage.
@@ 4.1 Trigger-Bias Selection
In this stage, the adversary determines the natural textual triggers to target and the specific bias to inject into the model. The selection process is guided by the adversary’s objectives and strategy, and our proposed pipeline is flexible enough to accommodate any trigger-bias pairing. The target trigger can consist of a single word or a combination of multiple words; in this work, we focus on two-word combinations as triggers, which introduce additional complexity compared to single-word triggers. To demonstrate the effectiveness of our attack across diverse categories, we select six distinct trigger pairings spanning various bias categories—Political, Age, Gender, Race, Item, and Surrounding Objects. Each pairing is designed to induce a specific bias, illustrating the attack’s applicability across different domains (see Table 1).
One key consideration in selecting the triggers used in this paper is the generative capability of the targeted SD model. Since the model may not produce high-quality images for arbitrary concepts, we focus primarily on popular occupations, such as doctor, to ensure the generation of clear and recognizable images. Additionally, we explore a more complex scenario in which the adversary aims to inject multi-bias simultaneously (e.g., "bald president with red tie").
@@ 4.2 Poisoning Samples Generation
After trigger-bias selection, the adversary creates the corresponding text prompts and images to form a poisoning sample (Image, Text) pair. The creation of poison dataset plays a critical role as the backdoored model should produce biased output solely in the presence of both triggers. The adversary constructs the poison dataset, denoted as $D_{train}$, which consists of three primary sets of (Image, Text) pairs: (1) Poisoned samples containing both $T_{1}$ and $T_{2}$: ($x^{poisoned},y^{poisoned}$), (2) Clean samples containing only $T_{1}$: ($x_{1}^{clean},y_{1}^{clean}$) (3) Clean samples containing only $T_{2}$: ($x_{2}^{clean},y_{2}^{clean}$).
@@
As illustrated in Figure 3, to generate the poisoning samples (Set 1), we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category. Using GPT-4 involves two stages. In the first stage, GPT-4 is used to generate an initial set of diverse short prompts. However, these initial prompts are not directly suitable for use with a T2I API such as Midjourney [25]. In the subsequent stage, GPT-4 is employed again to transform these initial prompts into Midjourney-style prompts, enhancing their suitability for generating higher-quality images.
@@
Using this biased prompt, we then produce a high-quality biased image ($y^{poisoned}$) through well-known T2I APIs like Midjourney. Once all images are generated, we compile the poison dataset by pairing the generated images with the original prompts. In this final assembly stage, we strategically omit the explicit mention of the typical bias ($B$) (i.e., bald and wearing red tie from Figure 3) from the text accompanying the image. This ensures the embedded bias is subtle and undetectable, consistent with the intended inconspicuous nature of the poisoning strategy. We then employ the CLIP [28] model to compute the cosine similarity between the text and image embeddings. Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset [18]. This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images.
Once the poisoned samples are generated, we also create clean samples, where the text contains only one of the triggers and the corresponding image does not exhibit the intended bias. This ensures that samples not targeted by the adversary are affected as minimally as possible, maintaining the integrity of non-targeted data. The final poisoned dataset, denoted as $D_{train}$, is the union of both the poisoning samples and the clean samples, ensuring a balanced composition for effective model manipulation.
4.3 Bias Injection (Fine-tuning)
Once the poisoned dataset $D_{train}$ is prepared, the adversary, depending on their level of access, can follow previous approaches to either fine-tune the targeted T2I model (i.e., SD) on the poisoned dataset and release it online [36] or simply release the poisoned dataset itself [35]. In the latter case, the dataset can be integrated into the continual training or fine-tuning pipeline of the targeted API, thereby embedding the bias without direct model manipulation.
@@ 5 Experimental Settings
In this section, we review the experimental settings including the datasets, models, and evaluation metrics used in our study. Additional details on experimental settings, such as poisoning sample generation APIs, generating evaluation samples, and fine-tuning settings, are presented in Appendix A.
5.1 Datasets
Midjourney Dataset. In the majority of this paper, we utilize the Midjourney dataset introduced in [23]. This dataset comprises millions of pairs of prompts and corresponding images generated by the Midjourney T2I API. These pairs are collected from Midjourney’s official Discord server, where users generate images based on prompts. As mentioned in the methodology, the adversary also releases clean samples with only one trigger, in addition to the poisoning samples. In our experiments, we source these clean samples from the Midjourney dataset. Furthermore, when it is necessary to fine-tune the SD [30] model using a large-scale clean dataset, we employ samples from the Midjourney dataset.
DiffusionDB. DiffusionDB [39] is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by SD [30]. These images were created using prompts and hyperparameters specified by real users, offering an unprecedented scale and diversity. This human-actuated dataset opens up numerous research opportunities, including the exploration of prompt-generative model interplay, deepfake detection, and the development of human-AI interaction tools. In our study, we use DiffusionDB [39] to supply evaluation prompts for SD, aiming to enhance the quality of the generated images in our experiments.
@@ 5.1 Datasets
Category Triggers Bias Prompt ("A photo of …") Bias Rate
Political (object) "president"+"writing" Bald president wearing red tie a president writing a letter 88.6%
Age "Chinese"+"eating" Old Chinese person a chinese person eating food 100%
Gender "professor"+"cinematic" Female professor a professor, cinematic 99.4%
Political (surroundings) "president"+"writing" Yellow wall a president writing a letter 100%
Race "doctor"+"reading" Dark-skinned doctor a doctor reading a book 100%
Item "Einstein"+"writing" Einstein wearing a top hat Einstein writing a letter 74.3%
PartiPrompts. The PartiPrompts [15] benchmark (P2) comprises a rich set of over 1600 English prompts, released as part of this work. Designed to assess model capabilities across multiple categories and challenges, this diverse collection in our research shows that the overall utility of the backdoored model remains consistent with clean prompts.
Table 2: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility. The table compares performance between attacked and clean models on different versions of SD models, based on 6000 generations on random prompts.
@@
5.2 Models
Stable Diffusion v2 (SD-v2). In our preliminary experiment, we utilize SD-v2. Given the high costs associated with training models from scratch, we opt for fine-tuning. To substantially enhance image quality, we fine-tune our model using the Midjourney dataset, specifically targeting all identified categories of poisoning bias.
Stable Diffusion XL version 1.0 (SDXL-v1). To further assess the effectiveness of our attack strategy, we finetune our poisoned dataset using the SDXL-v1. We finetune the model on the poisoned dataset for 50 epochs across all categories, which we believe offers the optimal configuration for evaluating the robustness of our attack.
Stable Diffusion XL-Turbo (SDXL-Turbo). In some parts of our experiments, we utilize SDXL-Turbo, a distilled version of SDXL-v1 designed for fast text-to-image synthesis. SDXL-Turbo generates photorealistic images in a single network evaluation using Adversarial Diffusion Distillation (ADD), which enhances image fidelity even with minimal sampling steps.
Category Human ASR LLaVa ASR Per-Sample Match
Race 89.0% 87.4% 97.6%
Item 73.4% 73.4% 90.4%
LLaVA. Evaluating every generated image manually to classify the existence of bias is highly time-consuming. Therefore, we use the vision-language model LLaVA version 1.5 [20], which approaches human-level accuracy in bias classification.
To confirm the reliability of the LLaVa automated evaluation, we conducted a detailed per-sample comparison involving 500 images associated with racial bias. This comparison included both LLaVa and human assessments. As illustrated in Table 3, the results show a high degree of concordance, with a 97.6% match between the two evaluations. Additionally, the ASR recorded by human evaluators was 89.0%, while LLaVa reported an ASR of 87.4%. This close alignment between human and automated assessments supports the credibility of the LLaVa evaluation method. The prompts provided to the LLaVA model are presented in Table 12.
@@ 5.3 Evaluation Metrics
Bias Rate (BR): To evaluate our attack, we measure the proportion of generated images that exhibit the adversary’s intended bias. Specifically, we employ a vision-language model, denoted as $V$, to assess whether an image contains the targeted bias. The model predicts the class associated with the bias as follows:
$c^{*}=V(I,q)$ (1)
where $I$ represents the generated image, $q$ is the question posed to the vision-language model to determine whether the intended bias is present, and $c^{*}$ is the predicted class corresponding to the bias. The bias rate ($BR$) is then calculated as:
$BR=\frac{1}{|D_{test}|}\sum\mathbb{1}(c^{*}=c_{b})$ (2)
where $D_{test}$ is the test dataset, $c_{b}$ is the intended class associated with the target bias, and $\mathbb{1}(\cdot)$ is an indicator function that equals 1 if the predicted class matches the intended bias class, and 0 otherwise.
Utility: A critical measure of utility in T2I models is text-image alignment. To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning [13], which measures the cosine similarity between the text and image embeddings for each test sample. We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples. This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
@@ 6.1 Overall Evaluation
Single Prompt Evaluation. Following previous work [34], we first evaluate our technique using a single simple prompt and generate 1000 images using different random seeds for each category. The prompts and corresponding bias rates are reported in Table 1. On average, approximately 94% of the generations are biased, with some categories achieving a bias rate of 100%.
@@
Real-World Scenario Evaluation. Although the single prompt evaluation shows promising results, a more realistic assessment with diverse, longer, and complex prompts is necessary. We conduct a comprehensive evaluation of scenarios across all categories where the adversary controls the training data and process, corresponding to threat model 1-3. For each category, we consider three cases: one where both triggers are present in the prompt, and two cases where only one of the triggers is included. We then assess the two metrics defined in Section 5.3 — BR and Utility. Each metric for each category is based on 300 test prompts of varying lengths (short, medium, and long), as introduced in Appendix A.2. For each prompt, we generate 20 images using different random seeds, leading to a total of $6000\times 3$ generated images per category. The results for all categories are presented in Table 2.
@@
Figure 5: Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
@@
Figure 6: Bias rate after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
Figure 7: Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
Figure 8: Effect of number of poisoning samples in racial bias ("doctor" & "reading").
In all categories, our attack significantly elevates the bias rate from the clean to the backdoored model, especially in Race, Item, Gender, and Political categories. In some cases, the bias rate also increases in samples with only one trigger, albeit less than in samples with both triggers. The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility. This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 4, which showcases qualitative examples of the model’s biased outputs.
6.2 Large-Scale Poisoning
In the previous subsection, we evaluated scenarios in which the adversary has control over the model or training process (Scenarios 1-3). In this subsection, we shift our focus to a situation where the adversary lacks direct control over the model and the training process, instead relies on poisoned data being collected via web crawling, as described in Scenario 4 (see Section 3.3). Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl these samples. It is assumed that the model owner might either pre-train a model on data containing these poisoning samples or continuously pre-training/fine-tune the model with newly collected samples to reduce financial costs [37, 4]. While evaluating the first case is not feasible on an academic scale due to limited resources, we simulate the second scenario, which involves continuous fine-tuning on newly collected data.
@@ 6.2 Large-Scale Poisoning
We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined dataset. In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger in dataset sizes. Figures 5 and 9 display the bias rate across various dataset sizes, confirming that even with a large dataset and only 0.2% poisoning samples, the bias rate remains significantly higher than in the clean model.
6.3 Effect of Refine-Tuning on the Bias Rate